Kinkaid Co. was incorporated at the beginning of this year and had a number of transactions. The following journal entries impacted its stockholders’ equity during its first year of operations.

General Journal Debit Credit
a. Cash 280,000
Common Stock, $25 Par Value 240,000
Paid-In Capital in Excess of Par Value, Common Stock 40,000
b. Organization Expenses 160,000
Common Stock, $25 Par Value 128,000
Paid-In Capital in Excess of Par Value, Common Stock 32,000
c. Cash 45,000
Accounts Receivable 19,000
Building 82,700
Notes Payable 59,600
Common Stock, $25 Par Value 57,100
Paid-In Capital in Excess of Par Value, Common Stock 30,000
d. Cash 123,000
Common Stock, $25 Par Value 80,000
Paid-In Capital in Excess of Par Value, Common Stock 43,000


Required:
2. How many shares of common stock are outstanding at year-end?
3. What is the total paid-in capital at year-end?
4. What is the book value per share of the common stock at year-end if total paid-in capital plus retained earnings equals $795,000?
Answer #1

Solution 2:

Outstanding common shares = total value of common stock / par value per share

= (240000+128000+57100+80000) / $25 = 20,204

Solution 3:

Total paid in capital = total common stock + total paid in capital in excess of par

= (240000+128000+57100+80000) + (40000+32000+30000+43000) = $650,100

Solution 4:

Book value per share = total paid in capital plus retained earnings / outstanding shares

= $795000/ 20204 = $39.35 per share